# OT-VP: Optimal Transport-guided Visual Prompting for Test-Time Adaptation

### v1

@@ Abstract
While Vision Transformers (ViTs) have demonstrated remarkable capabilities in learning representations, their performance is compromised when applied to unseen domains. Previous methods either engage in prompt learning during the training phase or modify model parameters at test time through entropy minimization. The former often overlooks unlabeled target data, while the latter doesn’t fully address domain shifts. In this work, our approach, Optimal Transport-guided Test-Time Visual Prompting (OT-VP), handles these problems by leveraging prompt learning at test time to align the target and source domains without accessing the training process or altering pre-trained model parameters. This method involves learning a universal visual prompt for the target domain by optimizing the Optimal Transport distance. With just four prompt tokens learned, OT-VP achieves a $5.0\%$ and $1.5\%$ increase in averaged accuracy across single-source and multi-source settings on three benchmark datasets, which is 1.2×1.2$\times$1.2 × and 1.5×1.5$\times$1.5 × the improvement of the state-of-the-art method, respectively.
@@ 1 Introduction
Figure 1: Motivation of our approach. (a) An ERM model trained on the source domain struggles to adapt to the target domain due to domain shifts. (b) Our method (OT-VP) optimizes a visual prompt by minimizing the Optimal Transport distance to align the target domain with the source domain without changing the decision boundary.
@@
Current approaches in prompt learning typically incorporate prompts during the training phase, leveraging source data to create representative prompts [49], or tuning prompts based on source domains at test time [10] for application in target domains. However, these conventional strategies often do not directly address the distribution shifts observed in the target domain, shown in Fig. 1 (a). Meanwhile, recent efforts focusing on prompt learning at test time [40, 27], are specifically designed for the CLIP (Contrastive Language-Image Pretraining) [37] model, targeting Vision-Language models exclusively. These methods generate input-specific prompts based on a limited set of augmented target data, potentially missing the broader differences between source and target domains. Adopting a universal prompt for the target domain leverages more data and reduces computation time. Yet, simply learning this universal prompt by minimizing entropy doesn’t effectively tackle domain shifts or enhance performance, as evidenced by our experiments in Sec. 4.3.
@@
To enable ViT models to adapt effectively at test time, it’s imperative to address the fundamental challenge of reducing the distribution gap between source and target domains. As shown in Fig. 1, this step is crucial for enhancing the model’s ability to generalize to new, unseen data. Many metrics, such as Maximum Mean Discrepancy (MMD), Kullback-Leibler (KL) divergence, and Optimal Transport (OT) distance, have been explored to address distribution shifts [3, 12, 20, 28]. Among these, Optimal Transport (OT) is distinguished as particularly effective due to its ability to utilize the geometry of the underlying space [35]. Numerous studies have demonstrated a strong correlation between OT distance and target performance, supported by thorough theoretical analyses [42, 29, 45]. Inspired by these insights, we propose Optimal Transport [17, 35] guided Test-Time Visual Prompting (OT-VP), a test-time distribution alignment method using prompt learning. Specifically, OT-VP explicitly aligns the Optimal Transport distance between the source representation, computed offline, and the target representation without accessing the training process and modifying the pre-trained model parameters. As illustrated in Fig. 2, for a given target dataset, we pass the unlabeled target image through the visual encoder with learnable prompts to get the target representation and pseudo-labels. Then we update the prompts by minimizing the OT distance. This process is repeated iteratively to ensure that the prompts are well-aligned with the target distribution.
@@
We comprehensively evaluate the efficacy of OT-VP across three standard benchmarks. Beyond the conventional multi-source DG setting—which leaves one domain out as the target while treating the remainder as sources—we also explore a single-source setting. The latter setting, often overlooked in existing literature [15, 49], involves using one domain as the source and a different one as the target. Training on a smaller set makes adaptation to the new target domain more challenging, providing a rigorous test of the model’s generalization ability. Based on ViT architecture, OT-VP achieves a stable and significant improvement in ERM, enhancing average accuracy across all three datasets by $5.0\%$ and $1.5\%$ in single-source and multi-source settings respectively. Notably, OT-VP surpasses state-of-the-art (SOTA) algorithms in both single-source and multi-source settings, underscoring its effectiveness and adaptability. Additionally, our method effectively reduces prediction entropy without directly minimizing it, thereby boosting model confidence and reliability on unseen data. Our contribution can be summarized as follows:
@@
  2. To the best of our knowledge, we are the first to learn universal prompts in the TTA setting. Our findings demonstrate that a universal prompt for a target domain can effectively align it with the source domain and enhance performance.
@@ 2 Related Work
Prompting for Vision Transformers. Vision Transformers (ViTs) have achieved state-of-the-art results in image classification [7, 25], yet adapting ViTs to unseen target domains without labels is still a challenge. Many recent methods enhance ViTs’ transferability by learning visual prompts as continuous, learnable vectors trained in an end-to-end manner with frozen model parameters [16, 2, 48]. However, these methods typically require labeled target data for prompt learning. For example, [49] introduced DoPrompt, a DG method that learns visual prompts for source domains and generates input-specific prompts at test time. However, its practical use is limited because it requires modifications to the training process. In contrast, our approach learns a universal prompt for the target domain, effectively minimizing the distribution shift between source and target domains without accessing the training setup.
Test-Time Adaptation (TTA). TTA [44, 47, 24] aims to improve the performance of pre-trained models when they are deployed on unseen domains. TTA strategies are typically bifurcated into two main categories based on their reliance on source data. The first category comprises approaches that adapt models to new domains without direct access to source data [22, 23, 36, 32]. Notable examples within this subset include Tent [47], which updates batch normalization parameters to minimize entropy. Similarly, T3A [15] innovates by creating and updating a set of dynamic prototypes, utilizing them to gauge the similarity between test samples and these prototypes, thereby facilitating adaptation based on feature resemblance.
Conversely, the second category, which our work aligns with, involves techniques that leverage source data to inform the adaptation process. Methods such as NORM [41] and DUA [30] adapt model normalization layers by aligning batch normalization statistics from the training phase with those observed in the target data. Similarly, ActMAD [31] and PromptAlign [40] update all model parameters or employ prompt-based strategies, respectively, utilizing a distribution matching loss that leverages knowledge of source domain statistics. This utilization of source data equips these methods with a foundational understanding of the distribution shift, thereby facilitating more effective adaptation to the target domain.
A significant portion of existing TTA literature centers around Convolutional Neural Networks (CNNs), whose methodologies are not directly transferrable to Vision Transformers (ViTs) due to architectural differences, notably the absence of batch normalization layers in ViTs [15]. This gap underscores the necessity for developing TTA strategies specifically tailored to the unique architecture of ViTs, which our work aims to address.
Innovative approaches like DePT[10], PromptAlign [40], and SwapPrompt [27] have introduced prompt-based strategies for TTA, learning input-specific prompts through test image augmentations. However, these methods encounter limitations due to the restricted number of augmentations available for each test image, struggling to minimize the distance between the comprehensive statistics of the source data and the augmented versions of a single test image. Our methodology circumvents this issue by optimizing a universal prompt across batches, which not only avoids the pitfalls of limited augmentations but also more effectively reduces the gap between source and target domains with only four prompt tokens.
@@ 3.1 Preliminaries
For any unlabeled target domain $\mathcal{D}^{t}$, TTA aims to adapt model $f$ or/and target input $x^{t}$ to bridge the performance gap under the assumption that the source domain and target domain share the same label set. In our approach, we employ a Vision Transformer as the model $f$, which remains fixed during adaptation.
@@ 3.4 Test-time Adaptation with OT-VP
The essence of OT-VP lies in calculating the OT distance between the source and target representations, underpinned by two distinct cost functions. The first, OT-VP-base (OT-VP-B), measures the cost between two representations devoid of label data, with the cost $c_{0}(\mathbf{z}^{s},\mathbf{z}^{t})$ defined as the Euclidean distance between source and target representations $\mathbf{z}^{s}:=h(\mathbf{x}^{s})$ and $\mathbf{z}^{t}:=h(\mathbf{x}^{s};\gamma)$:
@@
The second cost function, OT-VP, enriches this comparison by incorporating label or pseudo-label information, introducing a penalty term scaled by hyperparameter $\lambda$ for label mismatches between source and target data:
@@
where $\hat{y}^{t}:=f(\mathbf{x}^{t};\mathbf{\gamma}))$ represents the pseudo label derived from the pre-trained model using the adaptively learned prompts. Notably, when setting $\lambda$ to infinity in Eq. 7, the OT distance is the Wasserstein Distance [43], a well-established mathematical framework that has proven effective for measuring distances between distributions.
@@
Algorithm PACS VLCS OfficeHome Avg $\Delta$ Avg
ERM 64.5$\pm$3.2 63.8$\pm$2.2 66.7$\pm$0.7 65.0 0.0
DoPrompt 64.3$\pm$2.3 65.8$\pm$2.1 67.6$\pm$0.6 65.9 +0.9
Tent-C 64.4$\pm$2.4 65.0$\pm$3.2 66.6$\pm$1.2 65.3 +0.3
Tent-BN 69.0$\pm$2.9 58.5$\pm$2.3 67.6$\pm$0.9 65.0 0.0
T3A 71.2$\pm$2.8 68.3$\pm$2.2 68.1$\pm$0.8 69.2 +4.2
OT-VP-B 69.8$\pm$2.8 65.2$\pm$1.8 66.9$\pm$0.5 67.3 +2.3
OT-VP 73.5$\pm$2.7 68.4$\pm$2.3 68.1$\pm$0.7 70.0 +5.0
@@
Algorithm PACS VLCS OfficeHome Avg $\Delta$ Avg
ERM 87.0$\pm$0.1 78.5$\pm$0.3 73.6$\pm$0.1 79.7 0.0
DoPrompt 87.5$\pm$0.2 80.3$\pm$0.4 74.7$\pm$0.2 80.8 +1.1
Tent-C 87.2$\pm$0.2 78.8$\pm$0.5 73.6$\pm$0.2 79.9 +0.2
Tent-BN 87.2$\pm$0.2 78.5$\pm$0.3 74.6$\pm$0.3 80.1 +0.4
T3A 87.4$\pm$0.1 80.0$\pm$0.4 74.7$\pm$0.2 80.7 +1.0
OT-VP-B 87.3$\pm$0.2 80.2$\pm$0.3 74.3$\pm$0.3 80.6 +0.9
OT-VP 87.7$\pm$0.1 80.9$\pm$0.4 75.1$\pm$0.1 81.2 +1.5
@@ 4.1 Experimental Setup
Datasets. We evaluate OT-VP on three datasets for TTA: PACS [21], VLCS [8], and OfficeHome [46]. PACS [21] is composed of four domains: Photos, Art, Cartoon, and Sketch, containing 9,991 images in 7 classes. VLCS [8] comprises four real-world photographic datasets: VOC2007, LabelMe, Caltech, and SUN09, containing 10,729 images in 5 classes. OfficeHome [46] consists of four domains: Art, Clipart, Product, Real, containing 15,588 images in 65 classes.
@@
    Tent [47] is a TTA method that fine-tunes the batch normalization (BN) parameters to minimize prediction entropy during test time. Notably, ViTs do not contain BN layers. Following the methodology outlined by [15], we adapt Tent for use with ViTs in two distinct manners. Specifically, Tent-BN introduces a BN layer immediately preceding the linear classifier, allowing for the adjustment of normalization parameters within this BN layer. Tent-C modulates the entire classifier to reduce prediction entropy.
@@
Implementation details. Following [13], we partition the data from each domain into training and validation splits of $80\%$ and $20\%$, respectively, utilizing the larger split for training and the smaller one for model selection. For all experiments, we utilize the ViT-B16 model pre-trained on ImageNet, sourced from PyTorch [34]. Our training approach for ERM adheres to the hyperparameters specified by [49], incorporating a dropout rate of $0.1$ and a weight decay of $10^{-2}$. The learning rate is $5\times 10^{-6}$ for PACS and VLCS, and $10^{-5}$ for OfficeHome.
@@
For all baseline methods, we use their official implementation [^1] https://github.com/DequanWang/tent  [^2] https://github.com/matsuolab/T3A  [^3] https://github.com/zhengzangw/DoPrompt (see details in Appendix B.2). For implementation of our methods, we uses the DomainBed library [13]  [^4] https://github.com/facebookresearch/DomainBed. Adopting DoPrompt’s strategy, we use 4 prompts, optimized via the AdamW optimizer [26, 18] with a 0.1 learning rate on the target validation set. The target dataset’s batch size is 64, with prompt training capped at 5 epochs to avoid overfitting. We set $\lambda$ in Eq. 6 to $10^{8}$, as an estimate of infinity. All experiments were conducted using an RTX 6000 Ada GPU.
4.2 Results

### v2

@@ Abstract
Vision Transformers (ViTs) have demonstrated remarkable capabilities in learning representations, but their performance is compromised when applied to unseen domains. Previous methods either engage in prompt learning during the training phase or modify model parameters at test time through entropy minimization. The former often overlooks unlabeled target data, while the latter doesn’t fully address domain shifts. In this work, our approach, Optimal Transport-guided Test-Time Visual Prompting (OT-VP), handles these problems by leveraging prompt learning at test time to align the target and source domains without accessing the training process or altering pre-trained model parameters. This method involves learning a universal visual prompt for the target domain by optimizing the Optimal Transport distance.OT-VP, with only four learned prompt tokens, exceeds state-of-the-art performance across three stylistic datasets—PACS, VLCS, OfficeHome, and one corrupted dataset ImageNet-C. Additionally, OT-VP operates efficiently, both in terms of memory and computation, and is adaptable for extension to online settings. The code is available at https://github.com/zybeich/OT-VP.
@@ 1 Introduction
Figure 1: Motivation of our approach. (a) An ERM model trained on the source domain struggles to adapt to the target domain due to domain shifts. (b) Our method (OT-VP) optimizes a visual prompt by minimizing the Optimal Transport distance to align the target distribution (indicated as an ellipse) with the source distribution without changing the decision boundary.
@@
Current approaches in prompt learning typically incorporate prompts during the training phase, leveraging source data to create representative prompts [52], or tuning prompts based on labeled source data at test time [11] for application in target domains. However, these conventional strategies often do not directly address the distribution shifts observed in the target domain, shown in Fig. 1 (a). [52] generates input-specific prompts based on a limited set of augmented target data, potentially missing the broader differences between source and target domains. Adopting a universal prompt for the target domain leverages more data and reduces computation time. Yet, simply learning this universal prompt by minimizing entropy doesn’t effectively tackle domain shifts or enhance performance, as evidenced by our experiments in Sec. 4.4.
@@
To enable ViT models to adapt effectively at test time, it’s imperative to address the fundamental challenge of reducing the distribution gap between source and target domains. As shown in Fig. 1, this step is crucial for enhancing the model’s ability to generalize to new, unseen data. Many metrics, such as Maximum Mean Discrepancy (MMD), Kullback-Leibler (KL) divergence, and Optimal Transport (OT) distance, have been explored to address distribution shifts [3, 13, 22, 29]. Among these, Optimal Transport (OT) is distinguished as particularly effective due to its ability to utilize the geometry of the underlying space [18, 38]. Numerous studies have demonstrated a strong correlation between OT distance and target performance, supported by thorough theoretical analyses [44, 30, 47]. Inspired by these insights, we propose Optimal Transport-guided Test-Time Visual Prompting (OT-VP), a test-time distribution alignment method using prompt learning. Specifically, OT-VP explicitly minimizes the Optimal Transport distance between the source representation, computed offline, and the target representation without accessing the training process and modifying the pre-trained model parameters. As illustrated in Fig. 2, for a given target dataset, we pass the unlabeled target image through the visual encoder with learnable prompts to get the target representation and pseudo-labels. Then we update the prompts by minimizing the OT distance. This process is repeated iteratively to ensure that the prompts are well-aligned with the target distribution.
@@
We comprehensively evaluate the efficacy of OT-VP across three stylistic datasets: PACS, VLCS, and OfficeHome, as well as one corrupted benchmark: ImageNet-C. Beyond the conventional multi-source domain generalization (DG) setting, which typically leaves one domain out as the target while treating the remainder as sources, we also explore a single-source setting for stylistic datasets. This latter setting, often overlooked in existing literature [16, 52], involves using one domain as the source and a different one as the target. Training on a smaller dataset makes adaptation to the new target domain more challenging, thereby providing a rigorous test of the model’s generalization ability. Based on the ViT architecture, OT-VP achieves stable and significant improvements for the source pre-trained model, enhancing average accuracy across all three stylistic datasets by $5.0\%$ in the single-source setting and $1.5\%$ in the multi-source setting, respectively. Moreover, OT-VP shows an 11.5% improvement on the corrupted dataset. Notably, OT-VP surpasses state-of-the-art (SOTA) algorithms on both stylistic and corrupted datasets with fewer trainable parameters and less computational demand, underscoring its effectiveness and adaptability. Additionally, our method can be easily applied in more complex online settings. Our contribution can be summarized as follows:
@@
  2. Unlike existing methods, OT-VP does not require access to the training process or modifications to the pre-trained model’s parameters. Instead, OT-VP merely injects a few prompt tokens into the input layer, enhancing memory and computation efficiency and preventing catastrophic forgetting.
@@ 2 Related Work
Prompting for Vision Transformers. Vision Transformers (ViTs) have achieved state-of-the-art results in image classification [7, 27], yet adapting ViTs to unseen target domains without labels is still a challenge. Many recent methods enhance ViTs’ transferability by learning visual prompts as continuous, learnable vectors trained in an end-to-end manner with frozen model parameters [17, 2, 50]. These approaches, however, typically rely on labeled target data for prompt training. For instance, [52] introduces DoPrompt, a Domain Generalization (DG) method that learns visual prompts for source domains and generates input-specific prompts at test time, though it requires alterations to the training process, limiting its practical utility. Similarly, DePT [11] tunes prompts using labeled source data before adapting them to the target domain with a memory bank and a teacher-student system. In contrast, our approach learns a universal prompt for the target domain, effectively minimizing the distribution shift between source and target domains without requiring access to the training setup or any preliminary prompt tuning.
Test-Time Adaptation (TTA). TTA [46, 49, 24, 25, 26, 39, 34, 16, 32, 43, 33, 11, 36, 37, 31] aims to enhance the performance of pre-trained models when deployed in unseen domains. Traditionally, the majority of studies have focused on self-training and entropy regularization to align the model with the target data distribution. Notably, most existing TTA approaches have been developed for Convolutional Neural Networks (CNNs), and these methodologies are not directly applicable to Vision Transformers (ViTs) due to significant architectural differences, such as the absence of batch normalization layers in ViTs [16]. This disparity highlights the need for devising TTA strategies specifically tailored to the unique architecture of ViTs, a gap our work seeks to fill.
Furthermore, contemporary methods predominantly employ entropy minimization or consistency maximization often neglecting to directly confront domain shifts [49, 11]. These approaches typically involve modifications to the source pre-trained model’s parameters, which can lead to a risk of catastrophic forgetting [20, 10]. In contrast, our method uniquely leverages visual prompts while maintaining the entire pre-trained model frozen, thus effectively bridging the gap between source and target domains without altering underlying model parameters.
Algo. Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brit. Contr. Elastic Pixel. JPEG Avg. Gain
ERM 56.8 56.8 57.5 46.9 35.6 53.1 44.8 62.2 62.5 65.7 77.7 32.6 46.0 67.0 67.6 55.5 0.0
Tent-C 54.7 55.6 55.3 44.2 34.1 51.7 42.6 61.9 60.3 63.5 74.7 31.3 45.7 66.1 63.2 53.7 -1.8
Tent-BN 57.5 57.6 58.2 50.0 38.1 54.9 47.4 64.7 65.1 71.3 78.0 59.1 50.0 67.7 68.9 59.2 +3.7
Tent-LN 60.3 61.4 61.7 58.3 56.5 58.6 59.1 54.3 64.7 2.6 78.1 66.3 61.0 71.5 70.2 59.0 +3.5
T3A 56.4 56.9 57.3 47.9 37.8 54.3 46.9 63.6 60.8 68.5 78.1 38.3 50.0 67.6 69.1 56.9 +1.4
DePT 54.4 58.2 58.3 59.4 58.0 61.6 59.8 70.6 67.3 73.5 79.2 64.5 67.8 73.6 70.5 65.1 +9.6
OT-VP-B 60.9 63.2 64.0 56.8 51.7 59.1 56.0 68.1 63.4 73.3 80.0 68.3 62.9 72.7 72.8 64.9 +9.4
OT-VP 61.2 62.4 61.9 58.4 57.6 63.2 64.1 70.5 70.4 74.3 79.2 66.5 69.4 73.9 71.7 67.0 +11.5
DePT (online) 53.7 55.7 55.8 58.2 56.0 61.8 57.1 69.2 66.6 72.2 76.3 63.2 67.9 71.8 68.2 63.6 +8.1
OT-VP-B (online) 61.0 61.5 62.6 54.2 52.5 55.6 54.2 62.9 66.8 67.5 78.6 64.5 56.3 68.7 68.3 62.3 +6.8
OT-VP (online) 61.1 62.0 62.7 58.4 54.7 59.1 58.7 67.8 69.1 72.8 79.3 66.6 68.1 73.0 71.8 65.7 +10.2
Table 1: Results for ImageNet-C across 15 domains. OT-VP surpasses existing methods in both offline and online settings. The highest accuracy is highlighted in bold and the second highest is shown with an underline for clear distinction.
@@ 3.1 Preliminaries
For any unlabeled target domain $\mathcal{D}^{t}$, TTA aims to adapt model $f$ to bridge the performance gap under the assumption that the source domain and target domain share the same label set. In our approach, we employ a Vision Transformer as the model $f$, which remains fixed during adaptation.
@@ 3.4 Test-time Adaptation with OT-VP
The essence of OT-VP lies in calculating the OT distance between the target and pre-computed source representations, using two distinct cost functions. Note that the source representations are pre-computed prior to adaptation, and the source data is not utilized during test time. First, OT-VP-B (base) measures the cost between two representations devoid of label data, with the cost $c_{0}(\mathbf{z}^{s},\mathbf{z}^{t})$ defined as the Euclidean distance between source and target representations $\mathbf{z}^{s}:=h(\mathbf{x}^{s})$ and $\mathbf{z}^{t}:=h(\mathbf{x}^{s};\gamma)$:
@@
Second, OT-VP enriches this comparison by incorporating label or pseudo-label information, introducing a penalty term scaled by hyperparameter $\lambda$ for label mismatches between source and target data:
@@
where $\hat{y}^{t}:=f(\mathbf{x}^{t};\mathbf{\gamma}))$ represents the pseudo label derived from the pre-trained model using the adaptively learned prompts. Notably, when setting $\lambda$ to infinity in Eq. 7, the OT distance is the Wasserstein Distance [45], a well-established mathematical quantity that has proven effective for measuring distances between distributions.
@@
Algo. PACS VLCS OfficeHome Avg. Gain
ERM 64.5 63.8 66.7 65.0 0.0
DoPrompt 64.3 65.8 67.6 65.9 +0.9
Tent-C 64.4 65.0 66.6 65.3 +0.3
Tent-BN 69.0 58.5 67.6 65.0 0.0
T3A 71.2 67.7 68.1 69.0 +4.0
DePT 67.1 66.2 67.1 66.8 +1.8
OT-VP-B 69.8 65.2 66.9 67.3 +2.3
OT-VP 73.5 68.4 68.1 70.0 +5.0
@@
Algo. PACS VLCS OfficeHome Avg. Gain
ERM 87.0 78.5 73.6 79.7 0.0
DoPrompt 87.5 80.3 74.7 80.8 +1.1
Tent-C 87.2 78.8 73.6 79.9 +0.2
Tent-BN 87.2 78.5 74.6 80.1 +0.4
T3A 87.4 80.0 74.7 80.7 +1.0
DePT 87.3 80.1 74.8 80.7 + 1.0
OT-VP-B 87.3 80.2 74.3 80.6 +0.9
OT-VP 87.7 80.9 75.1 81.2 +1.5
@@ 4.1 Experimental Setup
Datasets. We evaluate our method on four benchmark datasets for TTA. Three stylistic datasets: PACS [23], VLCS [8], OfficeHome [48], and one corrupted dataset: ImageNet-C [15]. Details can be found in Appendix B.2
@@
    Tent [49] is a TTA method that fine-tunes the batch normalization (BN) parameters to minimize prediction entropy during test time. Notably, ViTs do not contain BN layers. Following the methodology outlined by [16], we adapt Tent for use with ViTs in two distinct manners. Specifically, Tent-BN introduces a BN layer immediately preceding the linear classifier, allowing for the adjustment of normalization parameters within this BN layer. Tent-C modulates the entire classifier to reduce prediction entropy. Additionally, following [37], Tent-LN directly adapts the layer normalization (LN) parameters.
@@
  4. 4.
    DePT [11] first tunes prompts on the source labeled data and then adapts the prompts to the target domain, utilizing an additional memory bank and a teacher-student system under a hierarchical self-supervised objective. Besides the visual prompts, DePT also updates the classifier parameters. In contrast, our method simplifies the approach by adding prompts only to the first transformer layer, without requiring any auxiliary models. We keep the entire pre-trained model frozen, making only the prompt tokens trainable.
@@
Implementation details. For all experiments, we utilize the ViT-Base model pre-trained on ImageNet-1k, sourced from timm [51]. For implementation of our methods, we use the DomainBed library [14]  [^1] https://github.com/facebookresearch/DomainBed. Adopting DoPrompt’s strategy, we use 4 prompts, optimized via the AdamW optimizer [28, 19] with a 0.1 learning rate on the target validation set. The target dataset’s batch size is 64, with prompt training capped at 100 and 50 training steps for ImageNet-C and the other three respectively to avoid overfitting. We set $\lambda$ in Eq. 7 to $10^{4}$. All experiments were conducted using an RTX 6000 Ada GPU.
4.2 Experiments on PACS, VLCS, and OfficeHome